# Stuart-Landau Oscillatory Graph Neural Network

### v1

@@ Abstract.
The code for SLGNN is available at [available upon publication]
@@ 1. Introduction
Graph neural networks (GNNs) are a class of neural networks designed to learn from graph-structured data by modeling the mutual influence between each node and its neighbors. These architectures have demonstrated effectiveness across diverse domains, ranging from social network analysis (fanGraphNeuralNetworks2019) to molecular property prediction (pmlr-v70-gilmer17a) and recommendation systems (pengSVDGCNSimplifiedGraph2022; damianouGraphFoundationModels2024).
However, classical GNN architectures, such as graph convolutional networks (GCN) (kipfSemiSupervisedClassificationGraph2016) and graph attention networks (GAT) (velickovicGraphAttentionNetworks2017), suffer from multiple pathologies stemming from their message-passing paradigm for aggregating features from neighboring nodes. These limitations include oversmoothing (oonoGraphNeuralNetworks2019; ntRevisitingGraphNeural2019; ruschSurveyOversmoothingGraph2023; zhang2025rethinking), overcorrelation (rothRankCollapseCauses2023), oversquashing (giovanni2024how; nguyenRevisitingOversmoothingOversquashing2022), and vanishing gradients (ruschGraphCoupledOscillatorNetworks2022).
To address these challenges, numerous alternative architectures have been proposed, such as Graph Neural ODEs that inherit properties from the physical systems on which they are based (poli2019graph; xhonneuxContinuousGraphNeural2019; chamberlainGRANDGraphNeural2021; Eliasof; thorpeGrandGraphNeural2022). Among them, a particularly promising class is the oscillatory graph neural networks (OGNN), based on coupled oscillator systems (nguyenCoupledOscillatorsGraph2023; ruschGraphCoupledOscillatorNetworks2022). In these models, each node feature is represented as an oscillator with coupling factors determined by the graph structure, drawing inspiration from models of electrical activity in biological neural networks These architectures have been shown to mitigate several pathologies, including vanishing and exploding gradients as well as oversmoothing (ruschGraphCoupledOscillatorNetworks2022; nguyenCoupledOscillatorsGraph2023).
@@ 2. Related Work
Continuous-depth GNN models based on graph ordinary differential equations (ODEs) have demonstrated significant success in circumventing known GNN limitations. Standard architectures like GATs can be interpreted as explicit discretizations of diffusion equations on graphs, and the corresponding continuous-time Graph Neural Diffusion (GRAND) model (chamberlainGRANDGraphNeural2021) has been shown to suffer substantially less from pathologies such as oversmoothing observed in discrete counterparts. Further improvements arise from augmenting the diffusion dynamics with reaction and advection terms, which can be rigorously proven to prevent convergence to node-wise constant steady states (thorpeGrandGraphNeural2022; GREAD_Choi; eliasof2024featuretransportation; GraphNeuralReactionDiffusionModels_Eliasof). Indeed, from the perspective of a dynamical system, the increasing similarity of the features in the long-time asymptotic is often a synonym of convergence of the trajectories of the different nodes toward a single steady state. This situation is directly linked to oversmoothing and always occurs when considering diffusion systems on a connected graph, in which case the unique steady state is the uniform one with magnitude depending on the initial condition. However, if the diffusion operator is learned in such a way as to almost disconnect the graph according to the task at hand, then convergence to the stable equilibrium can be significantly slowed. Additionally, adding reaction or advection terms to the model makes the steady state of the system dependent on the initial one, thus avoiding phenomena like oversmoothing and overcorrelation.
@@
A particularly promising class of continuous Neural ODEs consists of physics-inspired oscillatory models. Coupled Oscillatory Recurrent Neural Network (coRNN) (ruschCoupledOscillatoryRecurrent2020) employs coupled damped second-order harmonic oscillators to mitigate the vanishing and exploding gradient problems in residual neural networks for tackling long-sequence tasks. Unified Invertible Coupled Oscillatory RNN (UnICORNN) (ruschUnICORNNRecurrentModel2021) generalizes coRNN as a Hamiltonian system, improving memory efficiency while ensuring invertibility in time. Building on this, Linear Oscillatory State-Space models (LinOSS) (ruschOscillatoryStateSpaceModels2024) translates these dynamics from the standard RNN setting into a linear state-space model, enabling highly efficient computation via parallel scans rather than step-by-step recurrence. lanthalerNeuralOscillatorsAre2023 (lanthalerNeuralOscillatorsAre2023) further demonstrates the universality of the harmonic oscillatory networks for approximating underlying functions. Apart from harmonic-oscillator-based approaches, Artificial Kuramoto Oscillatory Neurons (AKOrN) (miyatoArtificialKuramotoOscillatory2024) propose a general-purpose oscillatory neuron as unit vectors evolving according to Kuramoto dynamics and with great adaptability with existing layer architectures such as convolutions and attention mechanisms.
@@
For the application of oscillatory models to graph tasks, the Graph Coupled Oscillator Network (GraphCON) (ruschGraphCoupledOscillatorNetworks2022) uses damped harmonic oscillators similar to those of coRNN, while Kuramoto Graph Neural Networks (KuramotoGNN) (nguyenCoupledOscillatorsGraph2023) model each feature as a fixed-amplitude oscillator with nonlinear coupling between oscillators. Both approaches provably reduce or eliminate oversmoothing as well as exploding and vanishing gradient problems. Notably, in coupled oscillator systems, a steady state is always the one where both phases and amplitudes are synchronized and possibly zero if damping factors are included. Therefore, in order to avoid oversmoothing/overcorrelation problems, oscillatory models admitting steady states other than the fully synchronized one are desirable. In this setting, our proposed Stuart-Landau architecture is of particular interest. Indeed, by enabling dynamic amplitude evolution alongside phase dynamics, it enriches the space of possible equilibrium states and enhances model expressiveness.
@@ 3.1. The Stuart-Landau Oscillator
Only two parameters play a role in the amplitude behavior, which gives us 4 main cases, (MPRT-SL; panteleyStabilityRobustnessStuartlandau2015; Reynolds_2025):
@@ 3.1.1. Coupled Stuart-Landau Oscillators
represents a stable synchronized asymptotic state for the system, (MPRT-SL; Reynolds_2025), represented in the final column of Table 1. However, depending on the connectivity matrix, other stable asymptotic states representative of incoherence (uniformly distributed around a smaller limit cycle) and other symmetric configurations can coexist with the fully synchronized state, (arenas-diaz-kurths-moreno-zhou), represented in the second to last column of Table 1.
On the other hand, in the subcritical case, the amplitudes converge to zero exponentially fast, regardless of the coupling, while the phase behavior can be characterized by (9) where each $\frac{r_{l}}{r_{j}}$ has converged to a finite value and thus can be absorbed into the network structure and each $r_{j}\to 0$ exponentially fast will no longer play a role in the phase dynamics (MPRT-SL). Thus one can see that the fully synchronized state $\dot{\phi}_{j}=\omega$ can be recovered in this regime, as well as the incoherent and other symmetric states found in (doi:10.1137/23M155400X).
@@ 3.2. Kuramoto oscillators
Careful observation of (9) would lead the reader to notice the similarity to the famous Kuramoto model of phase synchronization. In fact, Kuramoto derived his model from the system (7) via a phase reduction that removed the amplitude dependence of the oscillators (Kuramoto:1975ebm). The reduction, which is valid in the week coupling regime, involves taking the limits $\alpha,\beta\to\infty$ while keeping the ratio $\frac{\alpha}{\beta}$ fixed, forces each oscillator to take the amplitude $r_{j}=\sqrt{\alpha/\beta}$ so that the only relevant aspect of the model is the phase behavior. This is known as a phase reduction and gives rise to the model for identical Kuramoto oscillators (KO)
@@
The fully synchronized state represents the global minimum of the energy of the gradient flow and thus is always a steady state of the system and is exponentially stable. However, depending on the graph structure, other frequency synchronized states such that $\dot{\phi}_{j}=\dot{\phi}_{k}$ for all $j,k=1,\ldots,N$ while $\phi_{j}\neq\phi_{k}$ in general can also exist and be exponentially stable, e.g. for a ring structure (arenas-diaz-kurths-moreno-zhou). In (Taylor_2012) it was first shown that if a graph is dense enough, then the only stable outcome is the fully synchronized state, which is thus a global attractor, while in (doi:10.1137/23M155400X) it was shown that for a particular balanced graph structure, there are infinitely many distinct exponentially stable equilibria. Even further, under other specific graph structures there exist so-called twisted states and other balanced states which are not Lyapunov stable due to the existence of multiple zero eigenvalues, but nonetheless can be shown to be nonlinearly stable or neutrally stable (Sclosa_2024; 10.1063/1.2165594).
@@ 3.3.1. Coupled Harmonic Oscillators
For symmetric coupling $A$, it is classical, (Rao), to show that $x(t)=(x_{1}(t),\ldots,x_{n}(t))$ is given by
@@ 3.4.2. The coupled systems:
Active regimes: Starting with coupled undamped HOs, we see that the phase dynamics of each oscillator are entirely determined by the eigenvalues of the connectivity matrix $A$ and that the amplitudes are still determined by the initial data. Further, as seen in (ruschGraphCoupledOscillatorNetworks2022), the energy of the system is preserved. Therefore synchronization only occurs if the connectivity matrix has all identical eigenvalues, and each node has identical initial data, otherwise the behavior will in general be quasiperiodic, and nodes should remain distinguishable.
@@ 4. Start Landau Graph Neural Network
The implicit update of $R^{(l+1)}$ can be efficiently computed using either 1D Newton’s method or explicit inverse with Cardano’s formula, using the real root of the corresponding cubic equation. We discuss the details in Appendix D. When Newton’s method is used, it typically converges within 2 - 5 steps, depending on the set error tolerance. Both the encoder ENC and the the decoder DEC are single-layered MLPs. While ENC is complex-valued, from our experience the choice between a real- or complex-valued DEC does not noticeably affect model performance. Although this scheme requires repeated polar-cartesian conversion at each layer, empirically this introduces negligible computational overhead and does not impact model performance. This compares favorably with the alternative aforementioned 2D Newton’s method, which does not necessitate such conversions but is more computationally intensive.
@@ 5.1. Model Setups
  • GCN: The Graph Convolutional Network (kipfSemiSupervisedClassificationGraph2016) coupling is defined as $F_{\theta}^{(l)}(X^{(l)})=\sigma(D^{-\frac{1}{2}}AD^{-\frac{1}{2}}X^{(l)}W^{(l)})$, where $\sigma$ is LeakyReLU with tunable negative gradient, $A$ is the adjacency map of the graph augmented by self loops and $D$ is the diagonal matrix such that $D_{ii}=\sum_{j}A_{ij}$. $W^{(l)}$ is a learnable weight matrix.
  • GAT: The Graph Attention Network (velickovicGraphAttentionNetworks2017) coupling is defined as $F_{\theta}^{(l)}(X^{(l)})=\sigma(A^{(l)}(X^{(l)})X^{(l)}W^{(l)})$, where the adjacency matrix is computed with an attention mechanism. The entry for edge $(i,j)$ is given by:
@@
  • Tran: we also consider a feature-propagation-like setup similar to GRAND (chamberlainGRANDGraphNeural2021), where the feature update for each node is driven by a weighted sum of differences from its neighbors. Here the coupling function is defined as
@@
Second Order Harmonic Oscillator (GraphCON (ruschGraphCoupledOscillatorNetworks2022))
@@
The Kuramoto model treats the features of each node as the phase constrained on the surface of a hypersphere. The following Kuramoto architecture is proposed in (nguyenCoupledOscillatorsGraph2023)
@@
where $\omega$ is the natural frequency, $\mathrm{sin}(X_{k}^{(l)}-X_{i}^{(l)})$ the nonlinear coupling term, and $a_{\theta}^{(l)}$ represents a learnable attention mechanism. However, directly solving Equation (25) with explicit Euler is implementation-wise challenging, as it is difficult to constrain the phase features within the $0,2\pi]$ interval, leading to instability during backpropagation and significant underperformance. While prior work (nguyenCoupledOscillatorsGraph2023) used adaptive step-size solvers, these can be computationally expensive and prone to step-size underflow. To address this, as well as improving the adaptability of the model to different coupling functions, we propose an alternative numerical scheme by representing the phases as complex numbers on the unit circle $Z=e^{\mathbf{i}X}$, and rewrite the layer update as the following
@@ 5.2. Node Classification
We evaluate all models on transductive node classification tasks on both homophilic and heterophilic datasets, including Cora (mccallumAutomatingConstructionInternet2000), Citeseer (senCollectiveClassificationNetwork2008), Pubmed (namataQuerydrivenActiveSurveying2012), WebKB (peiGeomGCNGeometricGraph2020) (Texas, Wisconsin, Cornell), WikipediaNetwork (rozemberczkiMultiscaleAttributedNode2019) (Chameleon, Squirrel), heterophilous graphs (platonovCriticalLookEvaluation2023) (Amazon-rating). For homophilic graphs, we follow the data split schemes laid out by shchurPitfallsGraphNeural2018 (shchurPitfallsGraphNeural2018). As shown in Table 2 and 3, SLGNN variants consistently achieve competitive or state-of-the-art performance across both dataset types. Statistical significance tests are provided in Appendix B.

### v2

@@ Abstract.
The code and hyperparameters for SLGNN are available at https://github.com/kevvzhang/StuartLandauGNN
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792675isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Neural networksccs: Mathematics of computing Ordinary differential equationsccs: Mathematics of computing Graph algorithms
@@ 1. Introduction
Graph neural networks (GNNs) are a class of neural networks designed to learn from graph-structured data by modeling the mutual influence between each node and its neighbors. These architectures have demonstrated effectiveness across diverse domains, ranging from social network analysis (Fan et al., 2019) to molecular property prediction (Gilmer et al., 2017) and recommendation systems (Peng et al., 2022; Damianou et al., 2024).
However, classical GNN architectures, such as graph convolutional networks (GCN) (Kipf and Welling, 2016) and graph attention networks (GAT) (Veličković et al., 2017), suffer from multiple pathologies stemming from their message-passing paradigm for aggregating features from neighboring nodes. These limitations include oversmoothing (Oono and Suzuki, 2019; Nt and Maehara, 2019; Rusch et al., 2023; Zhang et al., 2025), overcorrelation (Roth and Liebig, 2023), oversquashing (Giovanni et al., 2024; Nguyen et al., 2022), and vanishing gradients (Rusch et al., 2022).
To address these challenges, numerous alternative architectures have been proposed, such as Graph Neural ODEs that inherit properties from the physical systems on which they are based (Poli et al., 2019; Xhonneux et al., 2019; Chamberlain et al., 2021; Eliasof et al., 2021; Thorpe et al., 2022). Among them, a particularly promising class is the oscillatory graph neural networks (OGNN), based on coupled oscillator systems (Nguyen et al., 2023; Rusch et al., 2022). In these models, each node feature is represented as an oscillator with coupling factors determined by the graph structure, drawing inspiration from models of electrical activity in biological neural networks. These architectures have been shown to mitigate several pathologies, including vanishing and exploding gradients as well as oversmoothing (Rusch et al., 2022; Nguyen et al., 2023).
@@ 2. Related Work
Continuous-depth GNN models based on graph ordinary differential equations (ODEs) have demonstrated significant success in circumventing known GNN limitations. Standard architectures like GATs can be interpreted as explicit discretizations of diffusion equations on graphs, and the corresponding continuous-time Graph Neural Diffusion (GRAND) model (Chamberlain et al., 2021) has been shown to suffer substantially less from pathologies such as oversmoothing observed in discrete counterparts. Further improvements arise from augmenting the diffusion dynamics with reaction and advection terms, which can be rigorously proven to prevent convergence to node-wise constant steady states (Thorpe et al., 2022; Choi et al., 2023; Eliasof et al., 2024a, b). Indeed, from the perspective of a dynamical system, the increasing similarity of the features in the long-time asymptotic is often a synonym of convergence of the trajectories of the different nodes toward a single steady state. This situation is directly linked to oversmoothing and always occurs when considering diffusion systems on a connected graph, in which case the unique steady state is the uniform one with magnitude depending on the initial condition. However, if the diffusion operator is learned in such a way as to almost disconnect the graph according to the task at hand, then convergence to the stable equilibrium can be significantly slowed. Additionally, adding reaction or advection terms to the model makes the steady state of the system dependent on the initial one, thus avoiding phenomena like oversmoothing and overcorrelation.
@@
A particularly promising class of continuous Neural ODEs consists of physics-inspired oscillatory models. Coupled Oscillatory Recurrent Neural Network (coRNN) (Rusch and Mishra, 2020) employs coupled damped second-order harmonic oscillators to mitigate the vanishing and exploding gradient problems in residual neural networks for tackling long-sequence tasks. Unified Invertible Coupled Oscillatory RNN (UnICORNN) (Rusch and Mishra, 2021) generalizes coRNN as a Hamiltonian system, improving memory efficiency while ensuring invertibility in time. Building on this, Linear Oscillatory State-Space models (LinOSS) (Rusch and Rus, 2024) translates these dynamics from the standard RNN setting into a linear state-space model, enabling highly efficient computation via parallel scans rather than step-by-step recurrence. Lanthaler et al. (Lanthaler et al., 2023) further demonstrates the universality of the harmonic oscillatory networks for approximating underlying functions. Apart from harmonic-oscillator-based approaches, Artificial Kuramoto Oscillatory Neurons (AKOrN) (Miyato et al., 2024) propose a general-purpose oscillatory neuron as unit vectors evolving according to Kuramoto dynamics and with great adaptability with existing layer architectures such as convolutions and attention mechanisms.
@@
For the application of oscillatory models to graph tasks, the Graph Coupled Oscillator Network (GraphCON) (Rusch et al., 2022) uses damped harmonic oscillators similar to those of coRNN, while Kuramoto Graph Neural Networks (KuramotoGNN) (Nguyen et al., 2023) model each feature as a fixed-amplitude oscillator with nonlinear coupling between oscillators. Both approaches provably reduce or eliminate oversmoothing as well as exploding and vanishing gradient problems. Notably, in coupled oscillator systems, a steady state is always the one where both phases and amplitudes are synchronized and possibly zero if damping factors are included. Therefore, in order to avoid oversmoothing/overcorrelation problems, oscillatory models admitting steady states other than the fully synchronized one are desirable. In this setting, our proposed Stuart-Landau architecture is of particular interest. Indeed, by enabling dynamic amplitude evolution alongside phase dynamics, it enriches the space of possible equilibrium states and enhances model expressiveness.
@@ 3.1. The Stuart-Landau Oscillator
Only two parameters play a role in the amplitude behavior, which gives us 4 main cases, (Millán et al., 2025; Panteley et al., 2015; Reynolds and Tudisco, 2025):
@@ 3.1.1. Coupled Stuart-Landau Oscillators
represents a stable synchronized asymptotic state for the system, (Millán et al., 2025; Reynolds and Tudisco, 2025), represented in the final column of Table 1. However, depending on the connectivity matrix, other stable asymptotic states representative of incoherence (uniformly distributed around a smaller limit cycle) and other symmetric configurations can coexist with the fully synchronized state, (Arenas et al., 2008), represented in the second to last column of Table 1.
On the other hand, in the subcritical case, the amplitudes converge to zero exponentially fast, regardless of the coupling, while the phase behavior can be characterized by (9) where each $\frac{r_{l}}{r_{j}}$ has converged to a finite value and thus can be absorbed into the network structure and each $r_{j}\to 0$ exponentially fast will no longer play a role in the phase dynamics (Millán et al., 2025). Thus one can see that the fully synchronized state $\dot{\phi}_{j}=\omega$ can be recovered in this regime, as well as the incoherent and other symmetric states found in (Sclosa, 2023).
@@ 3.2. Kuramoto oscillators
Careful observation of (9) would lead the reader to notice the similarity to the famous Kuramoto model of phase synchronization. In fact, Kuramoto derived his model from the system (7) via a phase reduction that removed the amplitude dependence of the oscillators (Kuramoto, 1975). The reduction, which is valid in the week coupling regime, involves taking the limits $\alpha,\beta\to\infty$ while keeping the ratio $\frac{\alpha}{\beta}$ fixed, forces each oscillator to take the amplitude $r_{j}=\sqrt{\alpha/\beta}$ so that the only relevant aspect of the model is the phase behavior. This is known as a phase reduction and gives rise to the model for identical Kuramoto oscillators (KO)
@@
The fully synchronized state represents the global minimum of the energy of the gradient flow and thus is always a steady state of the system and is exponentially stable. However, depending on the graph structure, other frequency synchronized states such that $\dot{\phi}_{j}=\dot{\phi}_{k}$ for all $j,k=1,\ldots,N$ while $\phi_{j}\neq\phi_{k}$ in general can also exist and be exponentially stable, e.g. for a ring structure (Arenas et al., 2008). In (Taylor, 2012) it was first shown that if a graph is dense enough, then the only stable outcome is the fully synchronized state, which is thus a global attractor, while in (Sclosa, 2023) it was shown that for a particular balanced graph structure, there are infinitely many distinct exponentially stable equilibria. Even further, under other specific graph structures there exist so-called twisted states and other balanced states which are not Lyapunov stable due to the existence of multiple zero eigenvalues, but nonetheless can be shown to be nonlinearly stable or neutrally stable (Sclosa, 2024; Wiley et al., 2006).
@@ 3.3.1. Coupled Harmonic Oscillators
For symmetric coupling $A$, it is classical, (Rao, 2004), to show that $x(t)=(x_{1}(t),\ldots,x_{n}(t))$ is given by
@@ 3.4.2. The coupled systems:
Active regimes: Starting with coupled undamped HOs, we see that the phase dynamics of each oscillator are entirely determined by the eigenvalues of the connectivity matrix $A$ and that the amplitudes are still determined by the initial data. Further, as seen in (Rusch et al., 2022), the energy of the system is preserved. Therefore synchronization only occurs if the connectivity matrix has all identical eigenvalues, and each node has identical initial data, otherwise the behavior will in general be quasiperiodic, and nodes should remain distinguishable.
@@ 4. Start Landau Graph Neural Network
The implicit update of $R^{(l+1)}$ can be efficiently computed using either 1D Newton’s method or explicit inverse with Cardano’s formula, using the real root of the corresponding cubic equation. We discuss the details in Appendix D. When Newton’s method is used, it typically converges within 2 - 5 steps, depending on the set error tolerance. Both the encoder ENC and the decoder DEC are single-layered MLPs. While ENC is complex-valued, from our experience the choice between a real- or complex-valued DEC does not noticeably affect model performance. Other readout types including modulus, phases, and phase differences do not improve and may sometimes degrade model performance. Although this scheme requires repeated polar-cartesian conversion at each layer, empirically this introduces negligible computational overhead and does not impact model performance. This compares favorably with the alternative aforementioned 2D Newton’s method, which does not necessitate such conversions but is more computationally intensive.
@@ 5.1. Model Setups
  • GCN: The Graph Convolutional Network (Kipf and Welling, 2016) coupling is defined as $F_{\theta}^{(l)}(X^{(l)})=\sigma(D^{-\frac{1}{2}}AD^{-\frac{1}{2}}X^{(l)}W^{(l)})$, where $\sigma$ is LeakyReLU with tunable negative gradient, $A$ is the adjacency map of the graph augmented by self loops and $D$ is the diagonal matrix such that $D_{ii}=\sum_{j}A_{ij}$. $W^{(l)}$ is a learnable weight matrix.
  • GAT: The Graph Attention Network (Veličković et al., 2017) coupling is defined as $F_{\theta}^{(l)}(X^{(l)})=\sigma(A^{(l)}(X^{(l)})X^{(l)}W^{(l)})$, where the adjacency matrix is computed with an attention mechanism. The entry for edge $(i,j)$ is given by:
@@
  • Tran: we also consider a feature-propagation-like setup similar to GRAND (Chamberlain et al., 2021), where the feature update for each node is driven by a weighted sum of differences from its neighbors. Here the coupling function is defined as
@@
Second Order Harmonic Oscillator (GraphCON (Rusch et al., 2022))
@@
The Kuramoto model treats the features of each node as the phase constrained on the surface of a hypersphere. The following Kuramoto architecture is proposed in (Nguyen et al., 2023)
@@
where $\omega$ is the natural frequency, $\mathrm{sin}(X_{k}^{(l)}-X_{i}^{(l)})$ the nonlinear coupling term, and $a_{\theta}^{(l)}$ represents a learnable attention mechanism. However, directly solving Equation (25) with explicit Euler is implementation-wise challenging, as it is difficult to constrain the phase features within the $0,2\pi]$ interval, leading to instability during backpropagation and significant underperformance. While prior work (Nguyen et al., 2023) used adaptive step-size solvers, these can be computationally expensive and prone to step-size underflow. To address this, as well as improving the adaptability of the model to different coupling functions, we propose an alternative numerical scheme by representing the phases as complex numbers on the unit circle $Z=e^{\mathbf{i}X}$, and rewrite the layer update as the following
@@ 5.2. Node Classification
We evaluate all models on transductive node classification tasks on both homophilic and heterophilic datasets, including Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008), Pubmed (Namata et al., 2012), WebKB (Pei et al., 2020) (Texas, Wisconsin, Cornell), WikipediaNetwork (Rozemberczki et al., 2019) (Chameleon, Squirrel), heterophilous graphs (Platonov et al., 2023) (Amazon-rating). For homophilic graphs, we follow the data split schemes laid out by Shchur et al. (Shchur et al., 2018). As shown in Table 2 and 3, SLGNN variants consistently achieve competitive or state-of-the-art performance across both dataset types. Statistical significance tests are provided in Appendix B.